OpenStudy (anonymous):

what is the radius of convergence for the series ((-1)^(n-1)*x^(3n)*2^n)/n

OpenStudy (anonymous):

is it |x|<1/cubicroot(2))?

OpenStudy (anonymous):

\[|x|<1/\sqrt[3]{2}\]

OpenStudy (anonymous):

Let's write that one in the formal form \[ a_n=\frac{(-1)^{n-1}x^{3n}2^n}{n} \] Using the ratio test, we have \[ \lim_{n \to \infty}\left| \frac{a_{n+1}}{a_n}\right| = \lim_{n \to \infty}\left| \frac{(-1)^{n}x^{3n+3}2^{n+1}}{n+1} : \frac{(-1)^{n-1}x^{3n}2^n}{n} \right|\] Therefore, \[ \lim_{n \to \infty}\left| \frac{a_{n+1}}{a_n}\right|=\lim_{n \to \infty}\left| \frac{2n x^3}{(n+1)}\right| =2|x|^3 \] Now, let \[ 2|x|^3 <1\] we have the radius of convergence is \[ |x| < \frac{1}{\sqrt[3]{2}} \]

OpenStudy (anonymous):

thanks just wanted to check that i did the right thing.

OpenStudy (anonymous):

yes, you're absolutely on the right track